OpenStudy (anonymous):

Graph:-3x^2+12y^2=84 Domain & Range Domain: all real numbers Range: {y/y<=-sqrt 7 or y >= sqrt 7}

OpenStudy (solomonzelman):

There are no restrictions for the domain. Range, seems to be as you said. all numbers, but nothing between positive and negative sqrt7. https://www.desmos.com/calculator/uqy5jcxkxm

OpenStudy (solomonzelman):

you seem to be right:)

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (solomonzelman):

Anytme.

OpenStudy (michele_laino):

please note that you graph is an hyperbola, whose equation is: \[\frac{ x ^{2} }{ (2\sqrt{7} )^{2}}-\frac{ y ^{2} }{(\sqrt{7})^{2}}=-1\]

OpenStudy (michele_laino):

@-EtherealEbullience

OpenStudy (anonymous):

Oh, thanks.

OpenStudy (michele_laino):

@-EtherealEbullience thanks!